# PF-LHM: 3D Animatable Avatar Reconstruction from Pose-free Articulated Human Images

### v1

@@ 3.6 Loss Function
To better capture complex clothing deformations, we adopt a diffused voxel skinning approach as proposed in [30, 20]. Given the predicted 3DGS parameters $\bm{\chi}=(\mathbf{p},\mathbf{r},\mathbf{f},\rho,\sigma)$, we transform the canonical avatar into target view space using voxel-based skinning.
@@
We render the animated Gaussian primitives via differentiable splatting to obtain an RGB image $\hat{I}$ and an alpha mask $\hat{M}$, based on the target camera parameters. Supervision is applied through the following photometric loss:
@@
where $\mathcal{L}_{\text{color}}$ and $\mathcal{L}_{\text{mask}}$ are L1 losses on RGB and alpha values respectively, and $\mathcal{L}_{\text{lpips}}$ is a perceptual loss measuring high-frequency feature similarity. We set the corresponding weights as $\lambda_{\text{rgb}}=1.0$, $\lambda_{\text{mask}}=0.5$, and $\lambda_{\text{per}}=1.0$.
@@
Empirically, we observe that using only mask supervision tends to encourage overly large Gaussian scales, especially near object boundaries, which leads to blurred renderings. To counteract this issue, we propose a Mask Distribution Loss $\mathcal{L}_{\text{dis}}$, which encourages uniform Gaussian distributions within human masks and sharper boundary representation. This is achieved by rendering an auxiliary mask $M_{\text{dis}}$ with fixed Gaussian parameters (opacity $\rho=0.95$, scale $\sigma=0.002$), and applying a L1 loss between $M_{\text{dis}}$ and the ground-truth human mask.
Furthermore, to reduce ambiguities in canonical space supervision, we adopt two additional geometric regularizers from LHM [31]: (1) the As Spherical As Possible loss $\mathcal{L}_{\text{ASAP}}$, which promotes isotropy in the 3D Gaussians, and (2) the As Close As Possible loss $\mathcal{L}_{\text{ACAP}}$, which preserves spatial coherence among neighboring primitives.
The combined geometric regularization term is defined as:
with empirically chosen weights: $\lambda_{\text{dis}}=0.5$, $\lambda_{\text{ASAP}}=20$, and $\lambda_{\text{ACAP}}=5$.
@@ 4 Experiments
Table 2: Comparison experiments with sparse-view input methods on public benchmark.
views InstantAvatar [14] GaussianAvatar [11] ExAvatar [23] PF-LHM-L
@@
2 19.324 0.929 0.136 87 s 13.572 0.863 0.180 90 s 25.534 0.953 0.049 8.5 m 26.544 0.962 0.026 1.65 s
4 20.860 0.928 0.129 143 s 18.201 0.916 0.112 134 s 26.563 0.956 0.043 15 m 26.979 0.962 0.025 2.23 s
8 20.833 0.926 0.134 254 s 21.830 0.942 0.073 216 s 27.573 0.960 0.040 32 m 27.311 0.963 0.023 4.11 s
16 21.075 0.930 0.119 451 s 23.461 0.949 0.059 376 s 28.108 0.962 0.039 1.22 h 27.888 0.965 0.022 10.64 s
Table 3: Comparison experiments with sparse-view input methods on causal videos.
views InstantAvatar [14] GaussianAvatar [11] ExAvatar [23] PF-LHM-L
2 22.624 0.950 0.079 87 s 18.455 0.935 0.089 90 s 26.723 0.969 0.031 8.5 m 28.229 0.974 0.016 1.65 s
4 23.627 0.954 0.073 143 s 21.073 0.949 0.063 134 s 27.532 0.970 0.028 15 m 28.450 0.974 0.016 2.23 s
8 23.548 0.954 0.071 254 s 23.980 0.960 0.043 216 s 28.431 0.972 0.026 32 m 28.643 0.975 0.015 4.11 s
16 23.420 0.955 0.065 451 s 24.943 0.962 0.037 376 s 28.916 0.973 0.024 1.22 h 28.924 0.977 0.014 10.64 s
Figure 3: Qualitative results of animatable human reconstruction from sparse-view inputs.
Implementation Details
We design three variants of our model with $N_{\text{layer}}=4,6,8$ layers of the PI-MT block, corresponding to PF-LHM-S (small), PF-LHM-M (medium), and PF-LHM-L (large), respectively. The models contain approximately 500 MB, 700 MB, and 1000 MB training parameters in total. We train the model by minimizing the training loss using the AdamW optimizer for 60,000 iterations. A cosine learning rate scheduler is employed, with a peak learning rate of 0.0001 and a warm-up period of 3,000 iterations. During each batch, we randomly sample a number of frames in the range of $[1,16]$ from a randomly selected training video. Input images are resized to have a maximum dimension of 1024 pixels. Training is performed on 32 A100 GPUs over five days. To ensure training stability, we apply gradient norm clipping with a threshold of 0.1. Additionally, we utilize bfloat16 precision and gradient checkpointing to enhance GPU memory and computational efficiency.
Training Dataset
For our model training, we utilize approximately 300,000 in-the-wild video sequences collected from public video repositories, along with over 5,173 3D public synthetic static human scans sourced from 2K2K [9], Human4DiT [35], and RenderPeople. Specifically, we employ a sampling ratio of 19:1 to draw training batches from the in-the-wild and synthetic datasets to balance generalization and view-consistency. To address view bias in the video data, we sample from a diverse range of perspectives as uniformly as possible, guided by the estimated global orientation of SMPL-X.
@@
We report PSNR, SSIM [43], and LPIPS [56] to assess rendering quality, and measure efficiency with GPU memory usage and inference time.
@@ 4.1 Comparison with Existing Methods
We conduct a comprehensive evaluation of PF-LHM by comparing it with three baseline methods for generating animatable human avatars from casually captured video sequences. We assess the efficiency and performance of our model using two types of datasets: one is a public benchmark that includes 20 video sequences from NeuMan [15], REC-MV [30], and Vid2Avatar [7], while the other comprises 24 casual video sequences collected via our smartphones.
Both Table 2 and Table 3 illustrate quantitative experiments evaluating our model against InstantAvatar [14], GaussianAvatar [11], and ExAvatar [23] on public and casual video sequences. Compared to the state-of-the-art (SOTA) baseline ExAvatar, our approach not only significantly accelerates the inference time but also yields comparable quantitative results. Specifically, for the model’s efficiency, our model promptly creates animable avatars in seconds while ExAvatar requires approximately 15 minutes to 1.2 hours, depending on the number of input images. In terms of model performance, SOTA methods typically require dozens of input images to achieve satisfactory metrics; Moreover, our model achieves more accurate results with substantially fewer inputs, and this capability improves as the number of input images increases. As shown in Figure 3, sparse input views lead to noticeable reconstruction artifacts with fitting-based frameworks, including geometric distortions and texture blurring. However, our PF-LHM achieves robust and high-fidelity reconstruction from sparse inputs and outperforms the SOTA baselines.
Table 4: Comparison with single-image method on pose animation using our in-the-wild dataset. * indicates we use 80000 query points in LHM for a fair comparison. + denotes the measured inference time of IDOL.
@@
AniGS [32] 1 17.023 0.858 0.087 15 minutes 24 GB
IDOL [60] 1 17.835 0.871 0.083 1.93 seconds + 23 GB
LHM-0.7B* [31] 1 20.338 0.920 0.047 4.77 seconds 21 GB
PF-LHM-M 1 20.335 0.919 0.047 1.11 seconds 11 GB
2 20.891 0.923 0.042 1.42 seconds 11 GB
4 21.172 0.925 0.039 2.02 seconds 13 GB
8 21.574 0.926 0.037 3.7 seconds 15 GB
16 21.752 0.928 0.037 9.15 seconds 18 GB

### v2

@@ 3.7 Loss Function
To better capture complex clothing deformations, we adopt a diffused voxel skinning approach as proposed in [26, 36]. Given the predicted 3DGS parameters, we transform the canonical avatar into target view space using voxel-based skinning.
@@
We render the animated Gaussian primitives via differentiable splatting to obtain an RGB image $\hat{I}$ and an alpha mask $\hat{M}$, based on the target camera parameters. Additionally, we generate the neural rendering results $\hat{I}_{\text{NR}}$ and $\hat{M}_{\text{NR}}$ using a lightweight DPT head. Supervision is applied through the following photometric loss:
@@
where $\mathcal{L}_{\text{color}}$ and $\mathcal{L}_{\text{mask}}$ represent L1 losses on the RGB and alpha values, corresponding to the 3D Gaussian splatting and neural rendering losses, respectively. The term $\mathcal{L}_{\text{lpips}}$ denotes a perceptual loss measures how similar the high-frequency features of neural-rendered and Gaussian Splat–rendered images are to those of the ground-truth images. The weights are set to $\lambda_{\text{rgb}}=1.0$, $\lambda_{\text{mask}}=0.5$, and $\lambda_{\text{per}}=1.0$.
@@
Although the photometric loss offers effective supervision in the target view space, the canonical representation remains under-constrained due to the ill-posed nature of monocular reconstruction. This limitation leads to deformation artifacts when warping the avatar into novel poses. To address this fundamental challenge, we employ Gaussian regularization loss $\mathcal{L}_{\text{reg}}$ to improve the learning of unobserved parts (see supplementary material for details).
@@ 4 Experiments
Table 2: Comparison with sparse-view input methods. The results are averaged on the public and smartphone-captured video sequences.
views InstantAvatar [17] GaussianAvatar [14] ExAvatar [29] LHM++
@@
2 23.337 0.944 0.091 3.6 min 22.008 0.947 0.067 3.8 min 26.588 0.965 0.036 8.5 min 27.835 0.970 0.018 1.08 sec
4 23.391 0.943 0.093 6.0 min 23.382 0.952 0.051 5.6 min 27.293 0.966 0.034 15 min 27.940 0.971 0.018 1.40 sec
8 23.650 0.943 0.090 10.6 min 24.106 0.956 0.048 9.0 min 28.006 0.968 0.031 32 min 28.147 0.971 0.017 2.09 sec
16 24.235 0.949 0.075 18.8 min 24.398 0.957 0.047 15.7 min 28.358 0.969 0.030 72 min 28.394 0.972 0.016 3.66 sec
64 25.271 0.956 0.055 54.0 min 25.308 0.960 0.040 66 min 29.587 0.971 0.027 264 min 28.422 0.972 0.016 17.94 sec
We design three variants of our model with $N_{\text{layer}}=4,6,8$ layers of the PI-MT block, corresponding to LHM++ (S), LHM++ (M), and LHM++ (L), respectively. The models comprise approximately 500M, 700M, and 900M training parameters in total. The training details can be found in the supplementary material. For brevity, LHM++ indicates LHM++ (M) unless otherwise specified.
@@
We report PSNR, SSIM [54], and LPIPS [70] to assess rendering quality. If not specifically written, and measure efficiency with GPU memory usage and inference time on a single NVIDIA A100-80G. For our evaluation benchmarks, we establish three datasets: public animation videos, smartphone-captured videos for sparse view reconstruction, and in-the-wild fashion videos for both single-view and sparse view settings. All evaluation datasets will be made publicly available for future research.
@@ 4.1 Comparison with Existing Methods
We first evaluate our method against three state-of-the-art fitting-based approaches designed for sparse inputs: InstantAvatar [17], GaussianAvatar [14], and ExAvatar [29]. For this evaluation, we utilize two datasets: (1) a public animation dataset comprising 23 video sequences from NeuMan [18], REC-MV [36], and Vid2Avatar [10]; and (2) a smartphone-captured video dataset containing 20 clips collected from in-the-wild environments.
Figure 3: Animatable human reconstruction comparisons from sparse images.
Table 3: Comparison with single-image methods on pose animation on our in-the-wild fashion video dataset.
@@
AniGS [38] 1 17.234 0.812 0.103 900.00 s 24.00 GB
IDOL [74] 1 17.912 0.847 0.097 1.93 s 23.12 GB
LHM-0.7B [37] 1 21.233 0.877 0.078 5.73 s 19.59 GB
8 21.761 0.883 0.073 24.76 s 25.42 GB
16 21.868 0.885 0.069 49.55 s 28.77 GB
64 21.852 0.885 0.069 381.48 s 51.73 GB
LHM++ 1 21.518 0.882 0.068 0.91 s 6.78 GB
8 22.208 0.886 0.060 2.09 s 7.35 GB
16 22.354 0.887 0.057 3.66 s 8.48 GB
64 22.363 0.887 0.057 17.94 s 29.51 GB
Table 2 summarizes quantitative experiments evaluating our model against baseline methods on public and smartphone-captured video datasets. Specifically, compared to ExAvatar, our method generates animatable avatars within seconds, whereas ExAvatar requires 15 minutes to over an hour. Additionally, unlike ExAvatar, which relies on dozens of input images, our model achieves higher accuracy with far fewer inputs. As illustrated in Figure 3, while LHM++ marginally outperforms ExAvatar with 16-view inputs, ExAvatar produces severe artifacts such as geometric distortions and texture blurring. In contrast, our method yields vivid and realistic animation results.
Figure 4 shows the avatar reconstruction results of LHM++ in canonical space from sparse images, demonstrating the robustness of the proposed model across diverse input categories, such as real and cartoon-style inputs.
Figure 4: Human reconstruction of LHM++ in canonical space from sparse image inputs.
Figure 5: Qualitative comparison with LHM. LHM++ matches LHM’s quality with a single input view and generates progressively more detailed results as view count increases. Please zoom in for better view.